Batteries are rated in terms of ampere-hours (A · h). For example, a battery that can deliver a current of 3.0 A for 5.0 h is rated at 15 A · h.

(a) What is the total energy, in kilowatt-hours, stored in a 12-V battery rated at 51 A · h? (kWh)

(b) At $0.12 per kilowatt-hour, what is the value of the electricity that can be produced by this battery? (¢)
Answer #1

(a) total energy = VI*t

= 12*51*3600 = 2203200 J

1 kWh = 3.6*10^6 J

= 2203200 / 3.6*10^6 = 0.612 kWh

(b) Value of electricity = 0.612*0.12 = 0.07344 $